OpenStudy (anonymous):

Under which conditions does the molar volume of a gas decrease? at 273 K and at 2.0 atm at 273 K and at 0.5 atm at 409.5 K and at 1.0 atm at 546 K and at 2.0 atm

OpenStudy (somy):

PV= nRT V= nRT/P ignore n and R so V= T/P

OpenStudy (somy):

check out all the values now

OpenStudy (anonymous):

136.5 546 409.5 273 Is it the first choice since it is the smallest?

OpenStudy (somy):

i guess thats how it is

OpenStudy (somy):

tho word 'decrease' kinda confuses me since how do we know if it decreased not if we don't know the original amount <.<

OpenStudy (anonymous):

Well thank you. :) Mind helping with a few more by chance?

OpenStudy (somy):

oh wait

OpenStudy (somy):

wait a sec try this one V/n= RT/P V/n= this is actually molar volume, i ignored n since its 1 but then i also ignored R try putting R values also

OpenStudy (somy):

i think if the result is less the 22.4 L that means it decreases

OpenStudy (anonymous):

That would make sense, and so that still leaves the first answer as correct.

OpenStudy (somy):

and answer is still first one lol

OpenStudy (somy):

yes :)

OpenStudy (anonymous):

You wouldn't happen to know how to solve this would you?: According to the equation 2Na + 2H2O mc012-1.jpg 2NaOH+H2, what mass of Na is required to yield 22.4 L of H2 at STP? (The atomic mass of Na is 22.99 u.) 1.00 g 2.00 g 23.0 g 46.0 g (I could close this question and open a new one if you prefer.)

OpenStudy (somy):

i think there was a picture?

OpenStudy (somy):

2Na + 2H2O --> 2NaOH+H2 i see

OpenStudy (somy):

well look there is formula volume= mole * 22.4 L so first thing lets find mole of H2

OpenStudy (somy):

volume is given as 22.4 , so put it into the formula and get the mole (don't get confused 22.4 in formula is STP condition, another 22.4L is given volume for H2)

OpenStudy (somy):

after you find a mole of H2 use this formula \[\frac{ mole~of~H_2 }{ coefficient~of~H_2 }=\frac{ mole~of ~Na }{ coefficient~of~Na }\]

OpenStudy (somy):

u will know mole of H2 and coefficient from reaction u need to find mole of Na and you know coefficient from reaction

OpenStudy (somy):

clear till now?

OpenStudy (anonymous):

Yeah sorry about that. I forget that the arrows don't copy correctly. and give me just a minute. I'm trying to let it all sink in. :)

OpenStudy (somy):

okay np

OpenStudy (somy):

sorry gtg i hope u get the upper part its simple ratio thing we've done before

OpenStudy (anonymous):

I've got it now. Thank you for your help. :)

OpenStudy (somy):

after this use mole= mass/Molecular mass so mass = mole*Molecular mass mole- u just found molar mass -given in the question thus find mass

OpenStudy (somy):

you are welcome :)